OpenStudy (anonymous):

How do you solve this,step by step : ln(x)*(2lnx+1)>=0

OpenStudy (zzr0ck3r):

\(ab\ge 0\) implies they are both positive or both negative. So let us assume they are both positive. \(\ln(x)\ge0\) when \(x\ge1\) and \(2\ln(x)+1\ge \) implies \(\ln(x)\ge -\frac{1}{2}\) implies \(x\ge \frac{1}{\sqrt{e}}\)\) Since \(1> \frac{1}{\sqrt{e}}\) we have that the whole thing is non-negative for \(x\ge 1\)

OpenStudy (zzr0ck3r):

Can you do the same thing for when they are both negative?

OpenStudy (zzr0ck3r):

Remembering the domain for \(\ln(x) \) is \(x>0\).

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (zzr0ck3r):

np